airplane problem

A small propeller airplane can comfortably achieve a high enough speed to take off on a runway that is 1/4 mile long. A large, fully loaded passenger jet has about thesame acceleration from rest, but it needs to achieve twice the speed to take off.What is the minimum runway length that will serve? Hint: You can solve this problemusing ratios without having any additional information
Answer #1
1/4 mile = 402.3 meters

vf^2 = vi^2 + 2*a*x

vi = 0

vf^2 = 2*a* 402.3 < Small plane


x = (vf^2 - vi^2) / (2a)

vi = 0

x = (vf^2)/(2a)

402.3 = (vf^2)/(2a)

NOW we have (2vf) for the big plane

??? = (2vf)^2 / (2a)

= 4vf^2 /(2a)
= 4 *(vf^2)/(2a)

= 4 *402.3 = 1 mile or 1609.2 meters <ANSWER

PLEASE RATE.
answered by: asdfg